# Strong order-one convergence of the Euler method for random ordinary differential equations driven by semi-martingale noises

### v7

@@ 5. Integral formula for the global pathwise error
$\tau^{N}(t)=\max_{t_{j}\leq t}\{t_{j}\}=\left[\frac{t}{\Delta t_{N}}\right]\Delta t_{N}=\left[\frac{tN}{T}\right]\frac{T}{N}.$ (5.2)
@@
$X_{t_{j}}-X_{t_{j}}^{N}=X_{t_{j-1}}-X_{t_{j-1}}^{N}+\int_{t_{j-1}}^{t_{j}}\left(f(s,X_{s},Y_{s})-f(t,X_{t}^{N},Y_{t})\right)\;\mathrm{d}s.$
@@
$=$ $\int_{t_{j-1}}^{t_{j}}\left(f(s,X_{s},Y_{s})-f(s,X_{t_{j-1}},Y_{s})\right)\;\mathrm{d}s$
@@
6. Basic estimates for the global error
Here we derive an estimate which is the basis for the convergence estimate. For that, we use the following classical discrete version of the Grownwall Lemma.
@@ 6. Basic estimates for the global error
$e_{j}\leq\sum_{i=0}^{j-1}a_{i}e_{i}+b,$ (6.1)
for every $j$, and where $a_{i},b\geq 0$. Then, $e_{j}\leq be^{\sum_{i=0}^{j-1}a_{i}},$ for all $j.$
Proof.
See e.g. [36, Lemma 1], with $m=j,$ $u_{m}=e_{j},$ $g_{m}=b\geq 0$ constant, and $K_{k}=a_{i}\geq 0.$
With that, we are ready to prove our basic estimate for the global error.
Proposition 6.1 (Basic strong estimate).
Under the 2.1, suppose further that Equation 2.6 holds. Then, for every $j=0,\ldots,N$,
@@
We estimate the first two integrals in Equation 5.1. For the first one, we use Equation 2.1 together with the assumed Equation 2.6, so that
@@
With these two estimates, we bound Equation 5.1 as
@@
This can be cast in the form of Equation 6.1, with a constant coefficient $a_{i}=c_{L}.$ Then, using the discrete Gronwall Lemma 6.1, we obtain
@@
for $j=1,\ldots,N$, where $\tau^{N}$ is given by Equation 5.2.
Taking the expectation of Equation 6.3, we obtain the desired result Equation 6.2. ∎
@@
Under the 2.1 and assuming conditions Equation 2.4, Equation 2.5 and Equation 2.6, we find that, for $j=1,\ldots,N,$
@@
Using that the paths of $\{X_{t}\}_{t}$ are almost surely absolutely continuous, adding and subtracting $f(\xi,0,Y_{\xi})$ to the integrand, and using assumption Item ii of the 2.1 along with Equation 2.6, we have almost surely that
@@
Integrating over $[0,t_{j}]$ and using Fubini’s theorem to exchange the order of integration,
@@
$=\Delta t_{N}\bigg{(}\int_{0}^{t_{j}}\mathbb{E}[\|f(s,0,Y_{s})\|]\;\mathrm{d}s$
@@ 7. Strong convergence for semi-martingale noises
Under the 2.1, suppose further that Equation 2.4, Equation 2.5 and Equation 2.6 hold; that $X_{0}^{N}$ is a first order approximation of $X_{0},$ in the sense that
@@
in $(t,x,y)\in I\times\mathbb{R}^{d}\times\mathbb{R}^{k}$, for suitable constants $c_{1},c_{2},c_{3},c_{4},c_{7}\geq 0$. Assume the semi-martingale noise $\{Y_{t}\}_{t\in I}$ is decomposed as $Y_{t}=Y_{0}+F_{t}+Z_{t}$, where $\{F_{t}\}_{t\in I}$ is an FV process, $\{Z_{t}\}_{t\in I}$ is a local martingale, and $F_{0}=Z_{0}=0,$ and assume they satisfy
@@
$\mathbb{E}\left[V(\{F_{t}\}_{t};I)\right]<\infty.$ (7.6)
@@
The expectation of the first term in Equation 7.8 is estimated using Equation 7.2, Equation 3.3, and Equation 7.1, so that
@@
where the first inner integral is a Lebesgue-Stieltjes integral and the second inner integral is a stochastic integral with respect to a local martingale. Since $\{F_{t}\}_{t\in I}$ is of finite variation, the first integral on the right hand side of Equation 7.10 is estimated using Equation 7.3, so that
@@
$\qquad\leq c_{4}\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\;\mathrm{d}s\;\|\mathrm{d}F_{\xi}\|\leq c_{4}\Delta t_{N}\int_{0}^{t_{j}}\;\|\mathrm{d}F_{\xi}\|\leq c_{4}\Delta t_{N}V(\{F_{t}\}_{t};[0,t_{j}]).$
@@
$\mathbb{E}\left[\left\|\int_{0}^{t_{j}}\int_{\tau^{N}(s)^{+}}^{s}D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}F_{\xi}\;\mathrm{d}s\right\|\right]\leq c_{4}\Delta t_{N}\mathbb{E}\left[V(\{F_{t}\}_{t};[0,t_{j}])\right].$ (7.11)
For the second integral on the right hand side of Equation 7.10, we once again switch the order of integration, then use the Itô isometry for local martingales and then again the bound Equation 7.3, so that
@@
$\leq\mathbb{E}\left[\left(\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}D_{y}f(\xi,X_{\tau^{N}(\xi)}^{N},Y_{\xi^{-}})\;\mathrm{d}s\;\mathrm{d}Z_{\xi}\right)^{2}\right]^{1/2}$
$\leq\mathbb{E}\left[\int_{0}^{t_{j}}\left(\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\|D_{y}f(\xi,X_{\tau^{N}(\xi)}^{N},Y_{\xi^{-}})\|\;\mathrm{d}s\right)^{2}\;\mathrm{d}[Z,Z]_{\xi}\right]^{1/2}$
$\leq\Delta t_{N}\mathbb{E}\left[\int_{0}^{t_{j}}\|D_{y}f(\xi,X_{\tau^{N}(\xi)}^{N},Y_{\xi^{-}})\|^{2}\;\mathrm{d}[Z,Z]_{\xi}\right]^{1/2}$
$\leq c_{4}\Delta t_{N}\mathbb{E}\left[\int_{0}^{t_{j}}\;\mathrm{d}[Z,Z]_{\xi}\right]^{1/2}$
$=c_{4}\Delta t_{N}\mathbb{E}\left[[Z,Z]_{t_{j}},\right]^{1/2},$
where we used that, for any $0\leq\xi\leq T,$ and for $\xi\leq s\leq\tau^{N}(\xi)+\Delta t_{N},$ we have $\tau^{N}(s)=\tau^{N}(\xi),$ so we can write $X_{\tau^{N}(s)}^{N}=X_{\tau^{N}(\xi)}^{N}.$ Now, using that $[Z,Z]_{0}=Z_{0}^{2}=0,$ we obtain
$\mathbb{E}\bigg{[}\bigg{\|}\int_{0}^{t_{j}}\int_{\tau^{N}(s)^{+}}^{s}D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}Z_{\xi}\;\mathrm{d}s\bigg{\|}\bigg{]}\leq c_{4}\Delta t_{N}\mathbb{E}\left[[Z,Z]_{t_{j}}\right]^{1/2}.$ (7.12)
@@
$\left\|f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi})-f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})-D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\Delta Y_{\xi}\right\|\leq\frac{c_{7}}{2}\left(\Delta Y_{\xi}\right)^{2}.$
@@
$\qquad\leq\frac{c_{7}}{2}\int_{0}^{t_{j}}\sum_{\tau^{N}(s)<\xi\leq s}\left(\Delta Y_{\xi}\right)^{2}\;\mathrm{d}s$
$\qquad\leq\frac{c_{7}}{2}\sum_{0<\xi\leq t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\left(\Delta Y_{\xi}\right)^{2}\;\mathrm{d}s$
$\qquad\leq\frac{c_{7}}{2}\Delta t_{N}\sum_{0<\xi\leq t_{j}}\left(\Delta Y_{\xi}\right)^{2}.$
@@
$\mathbb{E}\Bigg{[}\Bigg{\|}\int_{0}^{t_{j}}\sum_{\tau^{N}(s)<\xi\leq s}\bigg{(}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi})-f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\\ -D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\Delta Y_{\xi}\bigg{)}\;\mathrm{d}s\Bigg{\|}\Bigg{]}\leq\frac{c_{7}}{2}\Delta t_{N}\mathbb{E}\left[\left\|\sum_{0<\xi\leq t_{j}}\left(\Delta Y_{\xi}\right)^{2}\right\|\right].$ (7.13)
The final term in Equation 7.8 involves a Lebesgue-Stieltjes integral and is handled in a similar way, so that
@@
$\leq\frac{1}{2}\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\left\|D_{yy}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\right\|\;\mathrm{d}s\;\mathrm{d}[Y,Y]_{\xi}^{c}$
$\leq\frac{c_{7}}{2}\Delta t_{N}\int_{0}^{t_{j}}\;\mathrm{d}[Y,Y]_{\xi}^{c}\leq\frac{c_{7}}{2}\Delta t_{N}[Y,Y]_{t_{j}}^{c},$
where we used that $[Y,Y]_{0}^{c}=0.$ Taking the expectation,
$\mathbb{E}\left[\bigg{\|}\frac{1}{2}\int_{0}^{t_{j}}\int_{\tau^{N}(s)}^{s}D_{yy}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}[Y,Y]_{\xi}^{c}\;\mathrm{d}s\bigg{\|}\right]\leq\frac{c_{7}}{2}\Delta t_{N}\mathbb{E}\left[[Y,Y]_{t_{j}}^{c}\right].$ (7.14)
Putting the estimates Equation 7.9, Equation 7.10, Equation 7.11, Equation 7.12, Equation 7.13, and Equation 7.14 together and using that $0\leq[Y,Y]_{t}^{c}\leq[Y,Y]_{t}$ are increasing, we obtain, from Equation 7.8,
@@
$C_{T}=c_{1}T+c_{2}\sup_{0\leq s\leq T}\mathbb{E}\left[\|X_{\tau^{N}(s)}^{N}\|\right]+c_{3}\sup_{0\leq s\leq T}\mathbb{E}\left[\|Y_{s}\|\right]+c_{4}\mathbb{E}\left[V(\{F_{t}\}_{t};I)\right]\\ +c_{4}\mathbb{E}\left[[Z,Z]_{T}\right]^{1/2}+\frac{c_{7}}{2}\mathbb{E}\left[\left\|\sum_{0<s\leq T}\left(\Delta Y_{s}\right)^{2}\right\|\right]+\frac{c_{7}}{2}\mathbb{E}\left[[Y,Y]_{T}\right].$
Thanks to Equation 2.5, Equation 2.4, Equation 7.5, Equation 7.6, Equation 4.4, Equation 4.3, Equation 4.6, and Equation 4.7, the constant $C_{T}$ above is finite and independent of $N.$ Plugging Equation 7.15 into Equation 6.2 and applying Equation 6.4 and Equation 7.1 yield the desired strong order 1 estimate Equation 7.7. ∎
@@
8. Supplementary Material
In this supplementary material, we describe in details the numerical examples illustrating the main work. In all the examples driven by semi-martingale noises, we numerically verify the strong order 1 convergence of the Euler method. In the example with fractional Brownian motion, we verify the strong convergence of order $\min\{H+1/2,1\},$ where $H$ is the Hurst parameter. The order of convergence is displayed with a 95% confidence interval that perfectly matches the theory developed in the paper. They also show that the theoretical estimate is sharp.
One of the distinguishing powers of the Julia language is to work both as a high-level, interactive, scripting language (with its dynamically typed and just-in-time compilation system), and as a fast, high-performance language (with its ahead-of-time and aggressive compilation capabilities). It allows for both easy prototyping and the production of efficient end code. As such, it is growing quickly as a language suitable for scientific computing and computationally-demanding real-world applications.

### v8

@@ 5. Integral formula for the global pathwise error
$\tau^{N}(t)=\max_{t_{j}\leq t}\{t_{j}\}=\left\lfloor\frac{t}{\Delta t_{N}}\right\rfloor\Delta t_{N}=\left\lfloor\frac{tN}{T}\right\rfloor\frac{T}{N}.$ (5.2)
@@
$X_{t_{j}}-X_{t_{j}}^{N}=X_{t_{j-1}}-X_{t_{j-1}}^{N}+\int_{t_{j-1}}^{t_{j}}\left(f(s,X_{s},Y_{s})-f(t_{j-1},X_{t_{j-1}}^{N},Y_{t_{j-1}})\right)\;\mathrm{d}s.$
@@
$+\int_{t_{j-1}}^{t_{j}}\left(f(s,X_{s},Y_{s})-f(s,X_{t_{j-1}},Y_{s})\right)\;\mathrm{d}s$
@@
6. Basic bound for the global error
Here we derive a bound for the global error which is the basis for the convergence bound. For that, we use the following classical discrete version of the Gronwall Lemma, which can be deduced from e.g. [45, Lemma 1].
@@ 6. Basic bound for the global error
$e_{j}\leq\sum_{i=0}^{j-1}a_{i}e_{i}+b_{i},$ (6.1)
for every $j$, and where $a_{i},b_{i}\geq 0$, with $b_{i}$ nondecreasing. Then, for all $j,$ we have
$e_{j}\leq b_{j}e^{\sum_{i=0}^{j-1}a_{i}}.$ (6.2)
With that, we are ready to prove our basic bound for the global error.
Proposition 6.1 (Basic strong bound).
Under the 2.1, assume that Equation 2.6 holds. Then, for every $j=0,\ldots,N$,
@@
We bound the first two integrals in Equation 5.1. For the first one, we use Equation 2.1 together with the assumed Equation 2.6, so that
@@
Using these two bounds in Equation 5.1 yields
@@
This can be cast in the form of Equation 6.1, with a constant coefficient $a_{i}=c_{L}\Delta t.$ Then, using the discrete Gronwall Lemma 6.1, we obtain from (6.2) that
@@
for $j=1,\ldots,N$, where $\tau^{N}$ is given by Equation 5.2. Taking the expectation of Equation 6.4, we obtain the desired result Equation 6.3. ∎
@@
Under the 2.1, assume that Equation 2.4, Equation 2.5 and Equation 2.6 hold. Then, we find that, for $j=1,\ldots,N,$
@@
Using that the paths of $\{X_{t}\}_{t\in I}$ are almost surely absolutely continuous, adding and subtracting $f(\xi,0,Y_{\xi})$ to the integrand, and using assumption (ii) inside the 2.1 along with Equation 2.6, we have almost surely that
@@
Integrating over $0,t_{j}]$ and using Fubini’s theorem to exchange the order of integration,
@@
$\leq\Delta t_{N}\bigg(\int_{0}^{t_{j}}\mathbb{E}[\|f(s,0,Y_{s})\|]\;\mathrm{d}s$
@@ 7. Strong convergence for semi-martingale noises
Consider the Random ODE (1.1) and the Euler approximation (1.2), with the function $f=f(t,x,y),$ the initial condition $X_{0},$ and the noise $\{Y_{t}\}_{t\in I}$ satisfying the 2.1. Suppose that Equation 2.4, Equation 2.5 and Equation 2.6 hold; that $X_{0}^{N}$ is a first order approximation of $X_{0},$ in the sense that
@@
in $(t,x,y)\in I\times\mathbb{R}^{d}\times\mathbb{R}^{k}$, for suitable constants $c_{1},c_{2},c_{3},c_{4},c_{7}\geq 0$. Assume the semi-martingale noise $\{Y_{t}\}_{t\in I}$ is decomposed as $Y_{t}=Y_{0}+F_{t}+Z_{t}$, where $\{F_{t}\}_{t\in I}$ is an FV process, $\{Z_{t\wedge T}\}_{t\geq 0}$ is a martingale, and $F_{0}=Z_{0}=0,$ and assume they satisfy
@@
$\mathbb{E}\left[V(\{F_{t}\}_{t\in I};I)^{2}\right]<\infty,$ (7.6)
$\mathbb{E}\left[\|Z_{T}\|^{2}\right]<\infty.$ (7.7)
@@
The bound on the expectation of the first term in Equation 7.9 is obtained using Equation 7.2, so that
@@
where the first inner integral is a Lebesgue-Stieltjes integral and the second inner integral is a stochastic integral with respect to a local martingale. Since $\{F_{t}\}_{t\in I}$ is of finite variation, the bound on the first integral on the right hand side of Equation 7.11 is obtained using Equation 7.3 and switching the order of integration, so that
@@
$\qquad\leq c_{4}\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\;\mathrm{d}s\;\|\mathrm{d}F_{\xi}\|\leq c_{4}\Delta t_{N}\int_{0}^{t_{j}}\;\|\mathrm{d}F_{\xi}\|\leq c_{4}\Delta t_{N}V(\{F_{t}\}_{t\in I};[0,t_{j}]).$
@@
$\mathbb{E}\left[\left\|\int_{0}^{t_{j}}\int_{\tau^{N}(s)^{+}}^{s}D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}F_{\xi}\;\mathrm{d}s\right\|\right]\leq c_{4}\Delta t_{N}\mathbb{E}\left[V(\{F_{t}\}_{t\in I};[0,t_{j}])\right].$ (7.12)
For the second integral on the right hand side of Equation 7.11, we once again switch the order of integration, then use the Lyapunov inequality, the Itô isometry Equation 4.13 for local martingales, and then again the bound Equation 7.3, so that
@@
$\leq\mathbb{E}\left[\left\|\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}D_{y}f(\xi,X_{\tau^{N}(\xi)}^{N},Y_{\xi^{-}})\;\mathrm{d}s\;\mathrm{d}Z_{\xi}\right\|^{2}\right]^{1/2}$
$\leq c_{4}\Delta t_{N}\mathbb{E}\left[\sum_{l_{1},l_{2}=1}^{k}\int_{0}^{t_{j}}\left|\mathrm{d}[Z^{(l_{1})},Z^{(l_{2})}]_{\xi}\right|\right]^{1/2},$
where we used that, for any $0\leq\xi\leq T,$ and for $\xi\leq s\leq\tau^{N}(\xi)+\Delta t_{N},$ we have $\tau^{N}(s)=\tau^{N}(\xi),$ so we can write $X_{\tau^{N}(s)}^{N}=X_{\tau^{N}(\xi)}^{N}.$ Now, using that each quadratic variation and covariation is of finite variation, we obtain
$\mathbb{E}\bigg[\bigg\|\int_{0}^{t_{j}}\int_{\tau^{N}(s)^{+}}^{s}D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}Z_{\xi}\;\mathrm{d}s\bigg\|\bigg]\\ \leq c_{4}\Delta t_{N}\mathbb{E}\left[\sum_{l_{1},l_{2}=1}^{k}V(\{[Z^{(l_{1})},Z^{(l_{2})}]_{t}\}_{t\in I};[0,t_{j}])\right]^{1/2}.$ (7.13)
@@
$\left\|f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi})-f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})-D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\Delta Y_{\xi}\right\|\leq\frac{c_{7}}{2}\|\Delta Y_{\xi}\|^{2}.$
@@
$\qquad\leq\frac{c_{7}}{2}\int_{0}^{t_{j}}\sum_{\tau^{N}(s)<\xi\leq s}\left\|\Delta Y_{\xi}\right\|^{2}\;\mathrm{d}s$
$\qquad\leq\frac{c_{7}}{2}\sum_{0<\xi\leq t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\left\|\Delta Y_{\xi}\right\|^{2}\;\mathrm{d}s$
$\qquad\leq\frac{c_{7}}{2}\Delta t_{N}\sum_{0<\xi\leq t_{j}}\left\|\Delta Y_{\xi}\right\|^{2}.$
@@
$\mathbb{E}\Bigg[\Bigg\|\int_{0}^{t_{j}}\sum_{\tau^{N}(s)<\xi\leq s}\bigg(f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi})-f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\\ -D_{y}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\Delta Y_{\xi}\bigg)\;\mathrm{d}s\Bigg\|\Bigg]\leq\frac{c_{7}}{2}\Delta t_{N}\mathbb{E}\left[\left\|\sum_{0<\xi\leq t_{j}}\left\|\Delta Y_{\xi}\right\|^{2}\right\|\right].$ (7.14)
The final term in Equation 7.9 involves a Lebesgue-Stieltjes integral and is handled in a similar way, except that instead of the Itô isometry we just bound the Euclidian norm by the sum norm and obtain
@@
$=\bigg\|\frac{1}{2}\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}D_{yy}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}s\;\mathrm{d}[Y,Y]_{\xi}^{c}\bigg\|$
$\leq\frac{1}{2}\sum_{i=1}^{n}\left|\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\left(D_{yy}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}s\;\mathrm{d}[Y,Y]_{\xi}^{c}\right)^{(i)}\right|$
$\leq\frac{1}{2}\sum_{i=1}^{n}\left|\int_{0}^{t_{j}}\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\sum_{l_{1},l_{2}=1,\ldots,d}\partial_{y_{l_{1}}}\partial_{y_{l_{2}}}f_{i}(s,Y_{s^{-}})\;\mathrm{d}s\;\mathrm{d}[Y^{(l_{1})},Y^{(l_{2})}]_{\xi}^{c}\right|$
$\leq\frac{1}{2}\sum_{i=1}^{n}\sum_{l_{1},l_{2}=1,\ldots,d}\int_{0}^{t_{j}}\left(\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}\left|\partial_{y_{l_{1}}}\partial_{y_{l_{2}}}f_{i}(s,Y_{s^{-}})\right|\;\mathrm{d}s\right)\left|\;\mathrm{d}[Y^{(l_{1})},Y^{(l_{2})}]_{\xi}^{c}\right|$
$\leq\frac{1}{2}\sum_{i=1}^{n}\sum_{l_{1},l_{2}=1,\ldots,d}\int_{0}^{t_{j}}\left(\int_{\xi}^{\tau^{N}(\xi)+\Delta t_{N}}c_{7}\;\mathrm{d}s\right)\left|\;\mathrm{d}[Y^{(l_{1})},Y^{(l_{2})}]_{\xi}^{c}\right|$
$\leq\frac{c_{7}}{2}\Delta t_{N}\sum_{i=1}^{n}\sum_{l_{1},l_{2}=1,\ldots,d}V(\{[Y^{(l_{1})},Y^{(l_{2})}]_{t}\}_{t\in I};[0,t_{j}])$
Taking the expectation,
$\mathbb{E}\left[\bigg\|\frac{1}{2}\int_{0}^{t_{j}}\int_{\tau^{N}(s)}^{s}D_{yy}f(\xi,X_{\tau^{N}(s)}^{N},Y_{\xi^{-}})\;\mathrm{d}[Y,Y]_{\xi}^{c}\;\mathrm{d}s\bigg\|\right]\\ \leq\frac{c_{7}}{2}\Delta t_{N}\sum_{i=1}^{n}\sum_{l_{1},l_{2}=1,\ldots,d}\mathbb{E}\left[V(\{[Y^{(l_{1})},Y^{(l_{2})}]_{t}\}_{t\in I};[0,t_{j}])\right].$ (7.15)
Putting the bounds Equation 7.10, Equation 7.11, Equation 7.12, Equation 7.13, Equation 7.14, and Equation 7.15 together, we obtain, from Equation 7.9, that
@@
$C_{T}=c_{1}T+c_{2}T\sup_{0\leq s\leq T}\mathbb{E}\left[\|X_{\tau^{N}(s)}^{N}\|\right]+c_{3}T\sup_{0\leq s\leq T}\mathbb{E}\left[\|Y_{s}\|\right]\\ +c_{4}\mathbb{E}\left[V(\{F_{t}\}_{t\in I};I)\right]+c_{4}\mathbb{E}\left[\sum_{l_{1},l_{2}=1}^{k}V(\{[Z^{(l_{1})},Z^{(l_{2})}]_{t}\}_{t\in I};I)\right]^{1/2}\\ +\frac{c_{7}}{2}\mathbb{E}\left[\sum_{0<s\leq T}\left\|\Delta Y_{s}\right\|^{2}\right]+\frac{c_{7}}{2}\sum_{i=1}^{n}\sum_{l_{1},l_{2}=1,\ldots,d}V(\{[Y^{(l_{1})},Y^{(l_{2})}]_{t}\}_{t\in I};[0,t_{j}]).$
Thanks to Equation 2.4, Equation 2.5, Equation 2.6, Equation 3.3, Equation 7.5, Equation 7.6, Equation 7.7, Equation 4.4, Equation 4.3, Equation 4.9, Equation 4.7, and Equation 4.8, the constant $C_{T}$ above is finite and independent of $N.$ Plugging Equation 7.16 into Equation 6.3 and applying Equation 6.5 and Equation 7.1 yield the desired strong order 1 convergence Equation 7.8. ∎
@@
8. Numerical example
We have briefly mentioned, in the Introduction, several numerical examples confirming the strong order one convergence and therefore showing this rate is optimal. Here, we describe how the order of convergence is estimated numerically and then we give more details on one of the examples, namely the non-homogeneous linear example with different types of noises. The remaining examples can be found in the github repository [30].